OpenStudy (anonymous):

3. Ralph's age is one-third of his father Chuck's age. Ralph is 13. Write an equation to determine Chuck's age. A. 3c = 13 B. c+1/3=13 C. 1/3c=13 D. 1/3(13)=c

OpenStudy (ajprincess):

Let Chuck's age be c. Ralph"s age=1/3*c =13 That is 1/3c=13

OpenStudy (anonymous):

4. Using your answer from #3, find Chuck's age (c). A. c=4 and 1/3 B. c=12 and 2/3 C. c = 13 D. c = 39

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

hes 13?

OpenStudy (ajprincess):

Nope. Multiply by 3 both sides.

OpenStudy (ajprincess):

Ralph is only 13.

OpenStudy (anonymous):

D.39

OpenStudy (ajprincess):

ya u r right.

OpenStudy (anonymous):

thank you !

OpenStudy (ajprincess):

yw:)